# SOS-Match: Segmentation for Open-Set Robust Correspondence Search and Robot Localization in Unstructured Environments

### v2

@@
[…]
@@ I Introduction
A fundamental question to address in SLAM, as well as map-based localization approaches, is how to relate the current environment measurements, i.e., sensor inputs in the vicinity of the robot, efficiently and accurately to a reference map. We propose four requirements of robust correspondence search in the localization problem in unstructured environments. To resolve the ambiguities in the correspondence search of current observations and past observations, or across observations by different robots, the description should (1) provide high precision and recall. The method of description should (2) enable operation in an environment which lacks prominent landmarks, operating in a zero-shot approach i.e., without requiring significant engineering effort if the application domain changes. To allow operation over extended periods, it should (3) be robust to the variation in the appearance of the environment e.g., due to appearance change over seasonal time in the year. The description of the environment should (4) be modest in terms of memory use and computation time and communication bandwidth requirement for multi-agent scenarios.
To meet these requirements, we present SOS-Match, an open-set mapping and correspondence search pipeline that makes no prior assumptions about the content of the environment to extract and map objects from visually ambiguous unstructured settings, and uses only the geometric structure of the environment as cue for localization. Utilizing the segmentation model Segment Anything Model (SAM) [5] for front-end segment detection, our pipeline tracks detected segments across frames and prunes spurious detections to construct a map of consistent object masks, without requiring additional training per usage environment. We use a robust graph-theoretic data association method [6] to associate object locations within object maps, leveraging the geometric arrangement of landmarks and their relative position as cues for localization. Since many localization algorithms are expensive to run on platforms with limited computing resources, we formulate a windowed correspondence search that can trade off accuracy for computational cost. This is an especially suitable approach for drone localization over unstructured terrain, as we demonstrate through experiments with localization and loop closure detection in drone flights collected across different seasons.
@@ II-A Environment Representation
Descriptive and efficient environment representation lays the framework for robust localization. In visual SLAM, common environment representations include feature-based or object-based approaches. In feature-based SLAM, the environment is described by consistently detectable features within images that are distinct from one another like ORB [7], SIFT [8], SURF [9] or learned [10, 11] features. Several SLAM systems utilize these features for mapping, as shown in [12]. While feature-based SLAM is widely used, this approach poses a significant data handling challenge due to the substantial data volume it entails. In object-based SLAM methods, object detectors like YOLOv3 [13] can be used to extract a set vocabulary of objects in urban environments. Using object-based methods coupled with semantic labels can be advantageous in gathering contextual information about the environment while maintaining a compact map, as demonstrated in [14]. While the sparse nature of object-based mapping is promising for computational efficiency in large-scale operations, object classifiers are restricted to the types of objects that can be detected, limiting their use in unstructured environments.
@@ II-C Deep Learning in Visual Navigation
Some recent works in deep learning for visual navigation leverage foundational models [19], which are models trained in a self-supervised way that can accomplish many tasks without fine-tuning or additional training. Deep learning has been used to learn features in methods such as SuperPoint [11] and D2-Net [20]. In visual place recognition, learned global descriptors can facilitate robust scene recognition across viewpoints and differing illuminations. AnyLoc [21] is a technique for visual place recognition that uses DINOv2 [22], a self-supervised vision transformer model in combination with unsupervised feature aggregation using VLAD [23], which surpasses other VPR approaches in open-set environments, but degrades in cases where contents are similar across frames. Segment Anything uses deep learning to segment environments without prior assumptions or additional training [5].
@@ II-D Unstructured Environments
Unstructured environments are a challenging setting for many visual SLAM systems that assume urban-centric information such as the presence of lane markings or buildings. In unstructured environments, it is important to adequately extract meaningful features or recognize objects to aid in place recognition. Some approaches address road roughness and limited distinctive features in these settings by integrating a range of sensor modalities and strategies like using wheel odometry with visual tracking [24], integrating topological maps [25], and utilizing lidar point clouds [26]. These methods exhibit limitations tied to external hardware requirements, point cloud size constraints, susceptibility to structural changes, and reliance on the assumption of well-defined off-road trails.
@@ II-E Placement of This Work
SOS-Match utilizes a pre-trained foundation model for segmentation to construct object-based maps without any prior assumptions about the environment such as the presence of objects of specific classes. The generality of this framework allows it to localize successfully in unstructured environments with extreme illumination and structural changes while keeping map sizes compact enough to be shared between multiple agents.
@@ IV-A Dataset
We introduce the Båtvik seasonal dataset which includes six drone flights that travel a distance of approximately $3.5$ km over a coastal plot in southern Finland at an altitude approximately $100$ m over ground, each following the same trajectory plan as shown in Figure 1. We release this dataset for public use. The flights consist of drone images collected with a nadir-pointing camera as well as Inertial Measurement Unit (IMU) measurements, and we record autopilot output along with other telemetry data from an Ardupilot-based [33] drone flight controller. The flight takes place over an area that contains only a few buildings, and a large part of the trajectory takes place over a forest region, as well as above sea. This dataset represents flight of an UAV over a terrain that has naturally high ambiguity. We record this flight six times over many seasonal conditions, as illustrated in Figure 3 and outlined in Table I.
@@ IV-B Baseline approaches
Name
Icon
Time of flight
Description of appearance
Winter 1
2022-03-30 12:51
Snow coverage
Winter 2
2022-03-31 11:39
Snow coverage
Early Spring
2022-05-05 14:10
Some leaves
Late Spring
2022-05-25 12:33
Leaves in deciduous plants
Summer 1
2022-06-09 12:05
Full leaves, hard shadows
Summer 2
2022-06-09 12:28
Full leaves, hard shadows
In addition to feature-based approaches, we compare our method against a modern Visual Place Recognition (VPR) approach that uses global descriptors for images, AnyLoc [21]. AnyLoc outperforms universal place recognition pipelines NetVLAD [36], CosPlace [37] and MixVPR [38] in almost every evaluation, making it an appropriate benchmark that authors claim works across very different environmental and lighting conditions. In our implementation, we define a DINOv2 extractor following AnyLoc’s parameters at layer $31$ with facet value and $32$ clusters. We train a VLAD vocabulary of $32$ cluster centers on database images, generate global descriptors for each image in the query set, then compute the cosine similarity of the global descriptors of each image pair in each sequence.
@@ IV-D Precision and Recall
In Figure 4, we show precision and recall of each comparison method in the cross-seasonal localization case increasing visual discrepancy between flights from left to right cases.
@@
Our method provides better localization performance than the reference methods in the Summer 1 vs. Summer 2, Early Spring vs. Late Spring, and Late Spring vs. Summer 2 cases. Based on Figure 4, in the Winter 1 vs. Winter 2 cases, SuperPoint+SuperGlue appears to outperform our method. While there is a performance benefit in this case for SuperPoint+SuperGlue, in all cases, our method outperforms all reference methods with a significant margin in terms map size, and it outperforms all reference methods aside from AnyLoc in terms of seach time, as seen in Table IV-D. A low search time is a particularly critical characteristic for the description of a localization approach where search time is critical for the suitability for online implementation. Furthermore, a small map size is a key enabler of collaborative localization, where robots must share their local maps with limited network bandwidth.

### v3

@@
Annika Thomas1∗, Jouko Kinnari2∗, Parker C. Lusk1, Kota Kondo1, and Jonathan P. How1 Equal Contribution1A. Thomas, K. Kondo and J. How are with the Department of Aeronautics and Astronautics, Massachusetts Institute of Technology. {annikat, plusk, kkondo, jhow}@mit.edu.2J. Kinnari is with Saab Finland Oy, Salomonkatu 17B, 00100 Helsinki, Finland jouko.kinnari@saabgroup.com
@@ I Introduction
A fundamental question to address in SLAM, as well as map-based localization approaches, is how to relate the current environment measurements, i.e., sensor inputs in the vicinity of the robot, efficiently and accurately to a reference map. We propose four requirements of robust correspondence search in the localization problem in unstructured environments. To resolve the ambiguities in the correspondence search of current observations and past observations, or across observations by different robots, the description should (1) provide high precision and recall. The method of description should (2) enable operation in an environment which lacks prominent landmarks, operating in a zero-shot approach i.e., without requiring significant engineering effort if the application domain changes. To allow operation over extended periods, it should (3) be robust to the variation in the appearance of the environment e.g., due to appearance change over seasonal time in the year. The description of the environment should (4) be modest in terms of memory use, computation time and communication bandwidth requirement for multi-agent scenarios.
To meet these requirements, we present SOS-Match, an open-set mapping and correspondence search pipeline that makes no prior assumptions about the content of the environment to extract and map objects from visually ambiguous unstructured settings, and uses only the geometric structure of the environment as cue for localization. Utilizing the Segment Anything Model (SAM) [5] for front-end segment detection, our pipeline tracks detected segments across frames and prunes spurious detections to construct a map of consistent object masks, without requiring additional training per usage environment. We use a robust graph-theoretic data association method [6] to associate object locations within object maps, leveraging the geometric arrangement of landmarks and their relative position as cues for localization. Since many localization algorithms are expensive to run on platforms with limited computing resources, we formulate a windowed correspondence search that can trade off accuracy for computational cost. This is an especially suitable approach for drone localization over unstructured terrain, as we demonstrate through experiments with localization and loop closure detection in drone flights over time.
@@ II-A Environment Representation
Descriptive and efficient environment representation lays the framework for robust localization. In visual SLAM, common environment representations include feature-based or object-based approaches. In feature-based SLAM, the environment is described by consistently detectable features within a series of images like ORB [7], SIFT [8], SURF [9] or learned features [10, 11]. Several SLAM systems utilize these features for mapping, as shown in [12]. While feature-based SLAM is widely used, this approach poses a significant data handling challenge due to the substantial data volume it entails. In object-based SLAM methods, object detectors like YOLOv3 [13] can be used to extract a set vocabulary of objects in urban environments. Using object-based methods coupled with semantic labels can be advantageous in gathering contextual information about the environment while maintaining a compact map, as demonstrated in [14].
@@ II-C Deep Learning in Visual Navigation
Some recent works in deep learning for visual navigation leverage foundational models [19], which are models trained in a self-supervised way that can accomplish many tasks without fine-tuning or additional training. Deep learning has been used to learn features in methods such as SuperPoint [11] and D2-Net [20]. In visual place recognition, learned global descriptors can facilitate robust scene recognition across viewpoints and differing illuminations. AnyLoc [21] is a technique for visual place recognition that uses DINOv2 [22], a self-supervised vision transformer model in combination with unsupervised feature aggregation using VLAD [23], which surpasses other Visual Place Recognition (VPR) approaches in open-set environments, but degrades in cases where contents are similar across frames.
@@ II-D Unstructured Environments
Unstructured environments are a challenging setting for many visual SLAM systems that assume urban-centric information such as the presence of lane markings or buildings. Some approaches address road roughness and limited distinctive features in these settings by integrating a range of sensor modalities and strategies like using wheel odometry with visual tracking [24], integrating topological maps [25], and utilizing lidar point clouds [26]. These methods exhibit limitations tied to external hardware requirements, point cloud size constraints, susceptibility to structural changes, and reliance on the assumption of well-defined off-road trails.
@@ II-E Placement of This Work
SOS-Match utilizes a pre-trained foundation model for segmentation to construct object-based maps without any prior assumptions about the environment such as the presence of objects of specific classes. The generality of this framework allows it to localize successfully in unstructured environments with illumination and structural changes while keeping map sizes compact enough to be shared between multiple agents.
@@ IV-A Dataset
We introduce the Båtvik seasonal dataset which includes six $3.5$ km drone flights following the same trajectory plan as shown in Figure 1 at approximately $100$ m above ground, over many seasonal conditions as illustrated in Figure 3 and outlined in Table I. The flights consist of drone images collected with a nadir-pointing camera as well as Inertial Measurement Unit (IMU) measurements, and we record autopilot output along with other telemetry data from an Ardupilot-based [33] drone flight controller. The flight takes place over an area that contains only a few buildings, and a large part of the trajectory takes place over a forest region, as well as above sea. This dataset represents flight of an UAV over a terrain that has naturally high ambiguity.
@@ IV-B Baseline approaches
Name Icon Time of flight Description of appearance
Winter 1 [Uncaptioned image] 2022-03-30 12:51 Snow coverage
Winter 2 [Uncaptioned image] 2022-03-31 11:39 Snow coverage
Early Spring [Uncaptioned image] 2022-05-05 14:10 Some leaves
Late Spring [Uncaptioned image] 2022-05-25 12:33 Leaves in deciduous plants
Summer 1 [Uncaptioned image] 2022-06-09 12:05 Full leaves, hard shadows
Summer 2 [Uncaptioned image] 2022-06-09 12:28 Full leaves, hard shadows
In addition to feature-based approaches, we compare our method against a modern VPR approach that uses global descriptors for images, AnyLoc [21]. AnyLoc outperforms universal place recognition pipelines NetVLAD [36], CosPlace [37] and MixVPR [38] in almost every evaluation, making it an appropriate benchmark that authors claim works across very different environmental and lighting conditions. In our implementation, we define a DINOv2 extractor following AnyLoc’s parameters at layer $31$ with facet value and $32$ clusters. We train a VLAD vocabulary of $32$ cluster centers on database images, generate global descriptors for each image in the query set, then compute the cosine similarity of the global descriptors of each image pair in each sequence.
@@ IV-D Precision and Recall
In Figure 4, we show precision and recall of each comparison method localization case after time has passed, increasing visual discrepancy between flights from left to right cases.
@@
Our method provides better localization performance than the reference methods in the Summer 1 vs. Summer 2, Early Spring vs. Late Spring, and Late Spring vs. Summer 2 cases. Based on Figure 4, in the Winter 1 vs. Winter 2 cases, SuperPoint+SuperGlue appears to outperform our method. While there is a performance benefit in this case for SuperPoint+SuperGlue, in all cases, our method outperforms all reference methods with a significant margin in terms map size, and it outperforms all reference methods aside from AnyLoc in terms of search time, as seen in Table IV-D. A low search time is a particularly critical characteristic for the description of a localization approach where search time is critical for the suitability for online implementation. Furthermore, a small map size is a key enabler of collaborative localization, where robots must share their local maps with limited network bandwidth.
IV-E F1 Score: Performance Analysis by Viewpoint
To further analyze the performance of our method by viewpoint, we consider the cases in which an agent localizes within the map of another agent from the same viewpoint and when an agent passes over a place previously seen by another agent from a separate viewpoint. We calculate the average F1 score of different flights separated into the same viewpoint and different viewpoint, as shown in Figure 5.
Kuva 5: Average F1 scores of different cross-season cases. Bars indicate the performance from the same viewpoint (left) and from different viewpoints (right).
In this evaluation, we take the average F1 score calculated as the average value between when precision is tuned to at least $0.99$ and when recall is tuned to at least $0.99$. If recall cannot be tuned to at least $0.99$, we start at the highest value. Thus, these results demonstrate performance in cases that benefit from trading off between precision and recall.
In Figure 5, we see that all methods have a lower average F1 score in the different viewpoint setting than in the same viewpoint setting. We note that our method does not perform as well as ORB- and SIFT-based methods in the Summer 1 vs. Summer 2 case from multiple viewpoints due to our map representation being sparse and thus limited by field of view, but it maintains performance while others degrade when there are more visual variations between the flights.
In the same viewpoint case, Superpoint+Superglue surpasses our method in average F1 score. However, in the different viewpoint case, the Superpoint+Superglue fares unfavorably against our method and most comparison methods, suggesting that the approach is very sensitive to variation in viewpoint.
(a) Search time as function of window length.
(b) Search time as function of stride.
Kuva 6: Search time evaluations varying window length and stride.
(a) Varying window length, stride $=10$.
(b) Varying stride, window length $=50$.
Kuva 7: Precision and recall at different values of stride and window length (wl). Each graph is generated by varying minimum count of matches in window.
IV-F Ablation study
Due to the computational complexity of the correspondence search problem, we include tunable parameters for the windowed search approach that trade off between performance and runtime. Window length $\theta_{WL}$ is defined as the number of objects we consider at a time by their ID as assigned during map construction. Stride length $\theta_{SL}$ is the number of objects we skip as we slide the window along the entire traverse. In Figure 6, the search time throughout the entire traverse is evaluated. We demonstrate that increasing the window length increases runtime and increasing the stride length, and increasing the stride length decreases runtime.
To evaluate the performance of the system as the window length and stride length are tuned, we look at the precision and recall curves in Figure 7. Increasing the window length increases the performance of SOS-Match and decreasing the stride increases the performance of the system.
We find that the performance increases up to a window length of $50$ and a stride length of $10$, reflecting the parameters used in our experiments. These parameters enable tuning as a trade-off between performance and runtime that can be tailored to applications.
V Discussion
SOS-Match demonstrates the value of incorporating foundation models into front-end object detection and map construction in unstructured environments. Using segmentation in open-set unstructured settings such as dense forested regions provides sufficient geometric cues that are highly suitable for localization and loop closure detection. Our method also offers a significant speed improvement in search time and size reduction in map size in comparison to reference methods. We consider these major improvements towards satisfying the the requirements of a robust description of environment measurements for use in the localization problem, whose requirements we briefly listed in Sec. I.
We share the Båtvik seasonal dataset, which represents a challenging real-world scenario for visual navigation in unstructured environments, with significant ambiguities in the appearance of the environment. The data includes typical quality issues that occur in drones with hardware constraints such as image compression artifacts, which are useful for real-world evaluation. Our work reveals that most baseline methods are affected by even short time gaps between traverses, highlighting the need for robust visual approaches in these environments. The release of this dataset enables evaluation of robustness to changing seasons and visual conditions.
Our method does not fully account for uncertainty, and we plan to address cases with less favorable (non-bird’s eye view) triangulation geometry that may impact depth accuracy, as well as scenarios with significant odometry drift in future work.
The comparison of the localization performance against SuperPoint+SuperGlue suggests that it may be possible to find a balance between map size and localization capability by combining the information about the environment’s structure with a learning-based approach to description. We thereby plan to incorporate additional information about the objects, including semantic information in environments containing variable discernable objects, incorporating anisotropic object location covariance, and developing robust descriptors for geometry to enable faster correspondence search over large hypothesis spaces. Our proposed method uses size descriptors as a means for excluding putative matches where the size difference of objects is significant, and further search time reduction may be achievable if richer descriptors can be extracted for the segments.
VI Conclusion
We present SOS-Match, a framework for compact mapping and fast localization that is able to operate in open-set unstructured environments containing segmentable objects. The maps’ compactness supports a multi-agent scenario, facilitating efficient communication streams between agents. Experiments with the Båtvik seasonal dataset demonstrate the pipeline’s ability to align frames in a challenging unstructured environment with robustness to temporal and viewpoint variations. SOS-Match shows that segmentation provides geometric cues suitable for localization and robust correspondence search in unstructured environments.
Acknowledgement
This work was supported by Saab Finland Oy, Boeing Research & Technology, and the NSF GRFP under Grant No. 2141064. The dataset was collected as part of Business Finland project Multico (6575/31/2019). We thank the computational resources provided by the Aalto Science-IT project.
Viitteet
  • [1] R. Morales-Ferre, P. Richter, E. Falletti, A. de la Fuente, and E. S. Lohan, “A survey on coping with intentional interference in satellite navigation for manned and unmanned aircraft,” IEEE Communications Surveys & Tutorials, vol. 22, no. 1, pp. 249–291, 2020.
  • [2] J. Kinnari, R. Renzulli, F. Verdoja, and V. Kyrki, “Lsvl: Large-scale season-invariant visual localization for uavs,” Robotics and Autonomous Systems, vol. 168, p. 104497, 2023.
  • [3] C. Cadena, L. Carlone, H. Carrillo, Y. Latif, D. Scaramuzza, J. Neira, I. Reid, and J. J. Leonard, “Past, present, and future of simultaneous localization and mapping: Toward the robust-perception age,” IEEE Transactions on robotics, vol. 32, no. 6, pp. 1309–1332, 2016.
  • [4] Y. Tian, Y. Chang, F. Herrera Arias, C. Nieto-Granda, J. P. How, and L. Carlone, “Kimera-multi: Robust, distributed, dense metric-semantic slam for multi-robot systems,” IEEE Transactions on Robotics, vol. 38, no. 4, pp. 2022–2038, 2022.
  • [5] A. Kirillov, E. Mintun, N. Ravi, H. Mao, C. Rolland, L. Gustafson, T. Xiao, S. Whitehead, A. C. Berg, W.-Y. Lo et al., “Segment anything,” arXiv preprint arXiv:2304.02643, 2023.
  • [6] P. C. Lusk, K. Fathian, and J. P. How, “CLIPPER: A graph-theoretic framework for robust data association,” in 2021 IEEE International Conference on Robotics and Automation (ICRA), 2021, pp. 13 828–13 834.
  • [7] E. Rublee, V. Rabaud, K. Konolige, and G. Bradski, “ORB: An efficient alternative to SIFT or SURF,” in 2011 International conference on computer vision.   Ieee, 2011, pp. 2564–2571.
  • [8] D. G. Lowe, “Distinctive image features from scale-invariant keypoints,” International journal of computer vision, vol. 60, pp. 91–110, 2004.
  • [9] H. Bay, T. Tuytelaars, and L. Van Gool, “SURF: Speeded up robust features,” in Computer Vision–ECCV 2006: 9th European Conference on Computer Vision, Graz, Austria, May 7-13, 2006. Proceedings, Part I 9.   Springer, 2006, pp. 404–417.
  • [10] P.-E. Sarlin, D. DeTone, T. Malisiewicz, and A. Rabinovich, “Superglue: Learning feature matching with graph neural networks,” in Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, 2020, pp. 4938–4947.
  • [11] D. DeTone, T. Malisiewicz, and A. Rabinovich, “Superpoint: Self-supervised interest point detection and description,” in Proceedings of the IEEE conference on computer vision and pattern recognition workshops, 2018, pp. 224–236.
  • [12] I. Abaspur Kazerouni, L. Fitzgerald, G. Dooly, and D. Toal, “A survey of state-of-the-art on visual SLAM,” Expert Systems with Applications, vol. 205, p. 117734, 2022.
  • [13] J. Redmon and A. Farhadi, “YOLOv3: An incremental improvement,” arXiv preprint arXiv:1804.02767, 2018.
  • [14] J. Ankenbauer, P. C. Lusk, and J. P. How, “Global localization in unstructured environments using semantic object maps built from various viewpoints,” arXiv preprint arXiv:2303.04658, 2023.
  • [15] A. Radford, J. W. Kim, C. Hallacy, A. Ramesh, G. Goh, S. Agarwal, G. Sastry, A. Askell, P. Mishkin, J. Clark et al., “Learning transferable visual models from natural language supervision,” in International conference on machine learning.   PMLR, 2021, pp. 8748–8763.
  • [16] B. Douillard, J. Underwood, N. Kuntz, V. Vlaskine, A. Quadros, P. Morton, and A. Frenkel, “On the segmentation of 3D LIDAR point clouds,” in 2011 IEEE International Conference on Robotics and Automation.   IEEE, 2011, pp. 2798–2805.
  • [17] X. Zhao, W. Ding, Y. An, Y. Du, T. Yu, M. Li, M. Tang, and J. Wang, “Fast segment anything,” arXiv preprint arXiv:2306.12156, 2023.
  • [18] K. Kondo, C. T. Tewari, M. B. Peterson, A. Thomas, J. Kinnari, A. Tagliabue, and J. P. How, “PUMA: Fully decentralized uncertainty-aware multiagent trajectory planner with real-time image segmentation-based frame alignment,” arXiv preprint arXiv:2311.03655, 2023.
  • [19] R. Bommasani, D. A. Hudson, E. Adeli, R. Altman, S. Arora, S. von Arx, M. S. Bernstein, J. Bohg, A. Bosselut, E. Brunskill et al., “On the opportunities and risks of foundation models,” arXiv preprint arXiv:2108.07258, 2021.
  • [20] M. Dusmanu, I. Rocco, T. Pajdla, M. Pollefeys, J. Sivic, A. Torii, and T. Sattler, “D2-net: A trainable CNN for joint description and detection of local features,” in Proceedings of the ieee/cvf conference on computer vision and pattern recognition, 2019, pp. 8092–8101.
  • [21] N. Keetha, A. Mishra, J. Karhade, K. M. Jatavallabhula, S. Scherer, M. Krishna, and S. Garg, “AnyLoc: Towards universal visual place recognition,” arXiv preprint arXiv:2308.00688, 2023.
  • [22] M. Oquab, T. Darcet, T. Moutakanni, H. Vo, M. Szafraniec, V. Khalidov, P. Fernandez, D. Haziza, F. Massa, A. El-Nouby et al., “DINOv2: Learning robust visual features without supervision,” arXiv preprint arXiv:2304.07193, 2023.
  • [23] R. Arandjelovic and A. Zisserman, “All about VLAD,” in Proceedings of the IEEE conference on Computer Vision and Pattern Recognition, 2013, pp. 1578–1585.
  • [24] M. Grimes and Y. LeCun, “Efficient off-road localization using visually corrected odometry,” in 2009 IEEE International Conference on Robotics and Automation.   IEEE, 2009, pp. 2649–2654.
  • [25] T. Ort, L. Paull, and D. Rus, “Autonomous vehicle navigation in rural environments without detailed prior maps,” in 2018 IEEE international conference on robotics and automation (ICRA).   IEEE, 2018, pp. 2040–2047.
  • [26] R. Ren, H. Fu, H. Xue, X. Li, X. Hu, and M. Wu, “Lidar-based robust localization for field autonomous vehicles in off-road environments,” Journal of Field Robotics, vol. 38, no. 8, pp. 1059–1077, 2021.
  • [27] D. G. Lowe, “Distinctive image features from scale-invariant keypoints,” International Journal of Computer Vision, vol. 60, no. 2, pp. 91–110, 2004.
  • [28] R. I. Hartley and A. Zisserman, Multiple View Geometry in Computer Vision, 2nd ed.   Cambridge University Press, 2004.
  • [29] H. W. Kuhn, “The Hungarian method for the assignment problem,” Naval Research Logistics Quarterly, vol. 2, no. 1-2, pp. 83–97, 1955.
  • [30] F. Dellaert, M. Kaess et al., “Factor graphs for robot perception,” Foundations and Trends® in Robotics, vol. 6, pp. 1–139, 2017.
  • [31] F. Dellaert, “Factor graphs and GTSAM: A hands-on introduction,” Georgia Institute of Technology, Tech. Rep, vol. 2, p. 4, 2012.
  • [32] K. S. Arun, T. S. Huang, and S. D. Blostein, “Least-squares fitting of two 3-D point sets,” IEEE Transactions on Pattern Analysis and Machine Intelligence, vol. PAMI-9, no. 5, pp. 698–700, 1987.
  • [33] ArduPilot Community. (2022) Ardupilot - open source autopilot. Accessed: Dec 21, 2023. [Online]. Available: https://ardupilot.org
  • [34] M. A. Fischler and R. C. Bolles, “Random sample consensus: a paradigm for model fitting with applications to image analysis and automated cartography,” Communications of the ACM, vol. 24, no. 6, pp. 381–395, 1981.
  • [35] J. Sun, Z. Shen, Y. Wang, H. Bao, and X. Zhou, “LoFTR: Detector-free local feature matching with transformers,” CVPR, 2021.
  • [36] R. Arandjelovic, P. Gronat, A. Torii, T. Pajdla, and J. Sivic, “NetVLAD: CNN architecture for weakly supervised place recognition,” in Proceedings of the IEEE conference on computer vision and pattern recognition, 2016, pp. 5297–5307.
  • [37] G. Berton, C. Masone, and B. Caputo, “Rethinking visual geo-localization for large-scale applications,” in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, 2022, pp. 4878–4888.
  • [38] A. Ali-Bey, B. Chaib-Draa, and P. Giguere, “MixVPR: Feature mixing for visual place recognition,” in Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision, 2023, pp. 2998–3007.